12) In order to obtain 800 mL of a 2.5% lidocaine lotion. How much water must be used to dilute how much of a 10% lidocaine lotion? (Hint: water is 0% lidocaine lotion.)
Found 2 solutions by rothauserc, josgarithmetic:
Answer by rothauserc(4718)

let x be the amount of water and y be the amount of the 10% lidocaine lotion
:
x + y = 800 mL
:
0.025 * 800 = 20 mL of lidocaine
:
y = 800 -x
:
0.0 * x + 0.10 * (800 -x) = 20
:
80 -0.10x = 20
:
-0.10x = -60
:
x = 600
:
y = 800 -600 = 200
:
*************************************************************
mix 600 mL of water with 200 mL of the 10% lidocaine lotion
*************************************************************
:

Answer by josgarithmetic(39617)

That would be to dilute the 10% to one quarter of its strength. Does that help? Maybe less of Algebra and more of simple arithmetic.